OpenStudy (trisarahtops):

find dy/dx for 4 – xy = y^3.

OpenStudy (solomonzelman):

Can you differentiate y•x?

OpenStudy (trisarahtops):

that'd just be y

OpenStudy (solomonzelman):

You are using the product rule, and the chain rule every time you take derivative of y, and that is you multiply times y'.

OpenStudy (solomonzelman):

What would be the derivative of \(\color{#000000 }{ \displaystyle x\cdot f(x) }\) ?

OpenStudy (solomonzelman):

Use the product rule for that ...

OpenStudy (solomonzelman):

The derivative of \(f(x)\) is \(f'(x)\) and the derivative of \(x\) is?

OpenStudy (trisarahtops):

1

OpenStudy (solomonzelman):

for \(x\), and for f(x) ?

OpenStudy (trisarahtops):

no just x

OpenStudy (solomonzelman):

Wait, the derivative of \(x\) with respect to x is 1. Right?

OpenStudy (trisarahtops):

yes

OpenStudy (solomonzelman):

And the derivative of \(f(x)\) is \(f'(x)\) ok?

OpenStudy (trisarahtops):

right

OpenStudy (solomonzelman):

By the product rule (for differentiable functions of x, F and G) d/dx [F•G] = F' • G + F • G' Ok?

OpenStudy (trisarahtops):

yes

OpenStudy (solomonzelman):

d/dx [x•f(x)] = ? (you tell me)

OpenStudy (trisarahtops):

2f(x) ?

jimthompson5910 (jim_thompson5910):

h(x) = g(x) * f(x) h(x) = x * f(x) so, g(x) = x g ' (x) = 1 Using the product rule, we get h ' (x) = g ' (x) * f(x) + g(x)*f ' (x) h ' (x) = _________ (fill in the blank)

OpenStudy (solomonzelman):

You seem to struggle doing the product rule...

OpenStudy (solomonzelman):

Can you differentiate \(xe^x\) ? (Knowing that the derivative of \(e^x\) is \(e^x\) and derivative of x is 1)

OpenStudy (solomonzelman):

Product rule: d/dx (f•g) = f`g + fg`

OpenStudy (solomonzelman):

Want some examples of product rule?

OpenStudy (trisarahtops):

e^e ?

OpenStudy (solomonzelman):

correct yourself please.

OpenStudy (solomonzelman):

The derivative of \(e^x\) is just \(e^x\).

OpenStudy (trisarahtops):

okay so now what?

OpenStudy (solomonzelman):

Suppose I have the following function: \(\color{#000000 }{ \displaystyle h(x)=xe^x }\) Then the derivative of \(h(x)\) would be, \(\color{#000000 }{ \displaystyle \frac{d}{dx}~xe^x=\left(\frac{d}{dx} x \right) e^x+\left(\frac{d}{dx} e^x \right) x }\) The derivative of \(x\) (or \(x^1\)) is: \(\color{blue}{1}\cdot x^{\color{blue}{1}\color{red}{-1}}=1\cdot x^0=1\) (Knowing that \(y^0=1\) for all y besides y=0) The derivative of \(e^x\) is just \(e^x\). (If you like we can go through proving this via first principles of differentiation). So, we get; \(\color{#000000 }{ \displaystyle \frac{d}{dx}~(xe^x)=\left(\frac{d}{dx} x \right) e^x+\left(\frac{d}{dx} e^x \right) x }\) \(\color{#000000 }{ \displaystyle \frac{d}{dx}~xe^x=\left(1 \right) e^x+\left(e^x \right) x }\) \(\color{#000000 }{ \displaystyle \frac{d}{dx}~xe^x= e^x+\left(e^x \right) x }\)

OpenStudy (solomonzelman):

If you see something unfamiliar or questionable in what I said, then please ask.

OpenStudy (solomonzelman):

(take your time tho)

OpenStudy (trisarahtops):

I think I understand but how do I use this to solve the problem?

OpenStudy (mathmale):

Please note:\[4-xy=y^3\]

OpenStudy (mathmale):

is an implicit function. Theoretically, it could be solved for y, but that'd be hard and not worth the trouble.

OpenStudy (solomonzelman):

Did you get the product rule? Do you know the power rule (what I applied to differentiate the x)? Do you know the chain rule?

OpenStudy (mathmale):

So you must jump in and take the derivative of each term with respect to x.

OpenStudy (mathmale):

In other words, evaluate\[\frac{ d }{ dx}(4-xy=y^3)\]

OpenStudy (mathmale):

what is \[\frac{ d }{ dx}4?\]

OpenStudy (mathmale):

Hint: the derivative of a constant is always ___________

OpenStudy (trisarahtops):

zero

OpenStudy (mathmale):

Right. Now let's move on to

OpenStudy (mathmale):

\[\frac{ d }{ dx }[-xy]\]

OpenStudy (mathmale):

You could simplify matters by taking the neg. sign out.

OpenStudy (trisarahtops):

-y

OpenStudy (mathmale):

\[-\frac{ d }{ dx }xy\]

OpenStudy (trisarahtops):

or do you want y?

OpenStudy (mathmale):

solomon has gone over that with you already; please go back for a quick look at the product rule, and then at the derivative of xy with respect to x.

OpenStudy (mathmale):

Again: xy is a product, and we want the derivative of this product, so "do you want y?" is irrelevant.

OpenStudy (mathmale):

So, writing everything out formally, the derivative of xy is

OpenStudy (mathmale):

...I'll let Solomon type that for us.

OpenStudy (trisarahtops):

no i'm say the derivative of xy is y right?

OpenStudy (solomonzelman):

Looks like there is a little lack of prerequisite knowledge... but we can fix that if you bear with us.... it might take time, but wouldn't you agree it is worth it?

OpenStudy (solomonzelman):

Do you know the Power Rule? If you do please differentiate \(x^3\) for me.

OpenStudy (mathmale):

Trisa: Sorry, no. Solomon: Please stick to the derivative of xy first, then worry about the power rule later.

OpenStudy (mathmale):

Formally, the derivative with respect to x of xy is as follows:\[\frac{ d }{ dx }xy=x \frac{ dy }{ dx }+y \frac{ dx }{ dx }\]

OpenStudy (mathmale):

This is the applicaTION OF THE product rule for differentiation to xy. Note that we are assuming that y is a function of x. That's why I had to say NO to your question earlier.

OpenStudy (mathmale):

So, Tri-: can you simplify dx/dx?

OpenStudy (trisarahtops):

1

OpenStudy (mathmale):

Good! So, what do you have left?

OpenStudy (trisarahtops):

just y

OpenStudy (trisarahtops):

so now I simplify dy/dx?

OpenStudy (mathmale):

Yes.\[\frac{ d }{ dx }xy=x \frac{ dy }{ dx }+y \]

OpenStudy (trisarahtops):

d/dx xy = x y/2 +y

OpenStudy (mathmale):

No. Please move on to the last term, which is y^3. What is the derivative of that?

OpenStudy (mathmale):

I'm sorry, but x y/2 + y is not correct. Please compare it to my last previous equation, above.

OpenStudy (mathmale):

I'm getting the results I have by applying the product rule to xy.

OpenStudy (mathmale):

Try once more. The middle term is -xy. What is the derivative, with respect to x, of that?

OpenStudy (trisarahtops):

didn't u just tell me what the derivative of xy was just now??

OpenStudy (mathmale):

Yes. I want to be sure you agree with that before we move on.

OpenStudy (mathmale):

\[\frac{ d }{ dx }xy=x \frac{ dy }{ dx }+y\]

OpenStudy (trisarahtops):

x d/dx + y dx/dx

OpenStudy (mathmale):

Now enclose that in parentheses and write a negative sign in front. Then you'll have the derivative of -xy with respect to x, and assuming that y is a function of x.

OpenStudy (trisarahtops):

alright but how is dy/dx not simplified to y/x

OpenStudy (mathmale):

\[\frac{ d }{ dx }[-xy]=-(x \frac{ dy }{ dx }+y)\]

OpenStudy (mathmale):

I'm glad you asked that question. dy/dx cannot in any way be reduced. d is not a number, but an operator (command). dy/dx as a whole reads "the derivative of y with respect to x."

OpenStudy (trisarahtops):

ha nvm lets just continue

OpenStudy (trisarahtops):

so now I find the derivative of y^3?

OpenStudy (mathmale):

The last term we have to deal with is y^3. Find the derivative of that with respect to x. Be certain to apply the power rule AND the chain rule.

OpenStudy (trisarahtops):

3y^2

OpenStudy (mathmale):

Great, but you haven't yet applied the chain rule. Remember, y is assumed to be a function of x, and thus you must differentiate y with respect to x.

OpenStudy (mathmale):

So, the derivative of y^3 with respect to x is 3y^2 times what?

OpenStudy (trisarahtops):

sooo zero?

OpenStudy (mathmale):

The derivative of y with respect to x is denoted by dy/dx. You will see this again and again. Write the whole derivative of y^3 with respect to x here:

OpenStudy (mathmale):

Everything you need to do that is included in the most recent 3 entries, above.

OpenStudy (mathmale):

compare this to

OpenStudy (mathmale):

\[3y^2\frac{ dy }{ dx }\]

OpenStudy (mathmale):

sorry it's hard for you to understand this right now, but dy/dx is NOT equal to 1. Instead, dy/dx is a LABEL (like your name is a label) for "the derivative of y with respect to x."

OpenStudy (mathmale):

To get 3y^2, we used the power rule. We remember that y is assumed to be a function of x, so we follow the chain rule and write dy/dx immediately after 3y^2.

OpenStudy (mathmale):

In summary:\[\frac{ d }{ dx }4=0\]

OpenStudy (mathmale):

and ...

OpenStudy (mathmale):

\[-\frac{ d }{ dx }xy =-[x \frac{ dy }{ dx }+y]\]

OpenStudy (mathmale):

... and ...\[\frac{ d }{ dx }y^3=3y^2\frac{ dy }{ dx }\]

OpenStudy (mathmale):

I'm afraid you'll have to take my word for this for the time being. You can ask all the questions you want, perhaps tomorrow.

OpenStudy (mathmale):

So, put all 3 of these derivatives together to obtain the derivative of the implicit function 4-xy=y^3.

OpenStudy (mathmale):

Could you do that now, please? Don't bother with \[\frac{ d }{ dx}4,\]

OpenStudy (mathmale):

because it's zero. What's left?

OpenStudy (trisarahtops):

-[x dy/dx+ y] = 3y^2 dy/dx

OpenStudy (mathmale):

You are now writing an equation with the derivative of -xy on the left of the = sign and the deriv. of y^3 on the right. Yes, y ou have that correct.

OpenStudy (mathmale):

The last step of the process of finding the derivative is to SOLVE FOR dy/dx. Once again, dy/dx reads "the derivative of y with respect to x" and is not 1 and is not 0.

OpenStudy (mathmale):

There are 3 terms in the equation you've just written. 2 have dy/dx in them, and 1 does not. Can you agree with that?